OpenStudy (jenny.andrade):

Use the quadratic formula to solve the equation 5x^2 - 7x - 3 = 0 . Select one: a. x=(7+sqrt(109))/10 and x=(7-sqrt(109))/10 b. x=(7+sqrt(11))/10 and x=(7-sqrt(11))/10 c. No solution d. x=(-7+sqrt(109))/10 and x=(-7-sqrt(109))/10

OpenStudy (shamil98):

\[x = \frac{ -b \pm \sqrt{b^2 - 4(a)(c)} }{ 2a }\] Gather your values of a,b, and c. Using the standard form as a reference. \[ax^2 + bx + c = 0\] and then evaluate.

OpenStudy (jenny.andrade):

Soo 5x^2-7x+3=0

OpenStudy (jenny.andrade):

you add 5 and 2?